pls answer this simple question for me .... ^^ thanx The sum of three numbers is 4. The 1st, plus twice the second, plus the 3rd is 1. Three times the 1st, plus the second, minus the 3rd, is -2. Find the three numbers.
Answer by mananth(16946)

x+y+z= 4-------------- 1
x+2y+z =1-------------- 2
3x+2y-z =-2-------------- 3
consider equation 1 &2 Eliminate y
Multiply 1 by -2
Multiply 2 by 1
we get
-2x-2y-2z=-8
x+2y+z=1
Add the two
-x-z =-7 ------------- 4
consider equation 2 & 3 Eliminate y -
Multiply 2 by -1 x= 7
Multiply 3 by 1
we get
-x-2y-z =-1
3x+2y-z =-2
Add the two
2x-2z=-3 -------------5 5
Consider (4) & (5) Eliminate x
Multiply 4 by 2
Multiply (5) by 1
we get
-2x-2z= -14
2x-2z=-3
Add the two
-4z= -17
/-4
z=4.25

Plug the value of z in 5
2x-8.5=-3
2x=5.5
x=2.75
plug value of x & z in 1
2.75 +1y+4.25=4
1y=-2.75+-4.25+ 4
y=-3